ph of a solution obtained by mixing 100ml hcl soln of ph 2 and 400ml hcl soln of ph 3

Grade:12th pass

1 Answers

Bejoi Mathew
7 years ago
To solve this you have to follow this method,
As both solutions are acids you have to use ‘+’ to separate them.
solution A: 100ml and 10-2N
solution B: 400ml and 10-3N
N1V1 + N2V2 = NtVt
100* 10-2 + 400* 10-3 = Nt * 500
Solve for Nt,(i think you will be getting 28 * 10-4)
pH= -log (H+)
pH= – log (28 * 10-4)
PLEASE SOLVE FROM HERE ON.
HOPE YOU GRABBED THE CONCEPT.
BEST OF LUCK!